Sol Summary – April 22nd

Crew 280 Sol Summary Report 22-04-2023

Sol: 6

Summary title: Loud and Clean

Author’s name: Mariona Badenas-Agustí, Commander.

Mission status: Nominal

Sol activity summary:

The Hypatia I crew celebrated the start of the weekend with homemade pancakes and both savory and sweet combinations. After finishing breakfast, Carla Conejo González (Executive Officer & Biologist), Ariadna Farrés (Scientist & HSO), and Laia Ribas (GreenHab Officer) conducted an EVA around the base.

As the rest of the crew continued their scientific work indoors, the EVA team ventured outside to hone their drone flying skills and clear the area around the station, where there was a lot of space debris due to the high Martian winds of the past few days.

The team gathered for a well-deserved lunch at around 1PM. Today, Cesca Cufí ( Engineer) and Mariona Badenas-Agusti (Commander) served up a hot and hearty lunch of lentils infused with a tomato and vegetable sauce. This meal nourished the crew’s spirits and provided the perfect backdrop for a lively discussion about the future of the Hypatia Mars Association.

Later today, the crew will prepare dinner and play an exciting cardboard game, inspired by the very mission they are part of.

Look ahead plan: The crew is planning on holding a meditation session as well as a creative workshop dedicated to creating origami roses for St. George’s Day (also known as The Day of the Rose in Catalonia).

Anomalies in work: None.

Weather: Sunny. Low 42F/High 63F.

Crew physical status: Nominal.

EVA: We conducted an EVA around the base to fly a small drone and clean the terrain from space-debris.

Reports to be filed: Sol Summary, EVA request, EVA Report, Operations Report, GreenHab Report, Journalist Report, Pictures of the Day.

Support Requested: None.

Operations Report – April 22nd

Crew 280 Operations Report 22-04-2023
Name of person filing report: Neus Sabaté (Mission Specialist)
Non-nominal systems:
Robotic Observatory
Rover Spirit
Notes on non-nominal systems: Replacement parts of Robotic observatory are scheduled to be installed during crew 261 rotation. Corroded battery wire in Rover Spirit. Replacement expected for next week. Rover can only be used for short distance EVA.
Spirit rover used: no
Hours: 217.0
Beginning charge: –
Ending charge: –
Currently charging: yes
Opportunity rover used: no
Hours: 121.1
Beginning charge: –
Ending charge: –
Currently charging: yes
Curiosity rover used: no
Hours: 230.1
Beginning charge: –
Ending charge: –
Currently charging: yes
Perseverance rover used: no
Hours: 261.5
Beginning charge: –
Ending charge:-
Currently charging: yes
General notes on rovers: last check made on SOL 5 (12:05 PM)
Summary of Hab operations: put notes here
WATER USE: 31 gal (from 6PM Sol4 to 6PM Sol5)
Water (static tank): 366 gal (at 6:00PM)
Static tank pipe heater: off
Static tank heater: off
Toilet tank emptied: yes
Summary of internet: used from 18:00 to 21:00 for personal communication and reports.
Summary of suits and radios:
Suits: 1, 4 and 7
Radios: nominal
All radio and spacesuit batteries checked before EVA: Yes. All nominal.
Summary of GreenHab operations: put notes here
WATER USE: 14 gallons
Heater: Off
Supplemental light: Off
Harvest: 234 g cucumber, 112 g tomatoes, 4 g chives, 15 g jazz mix, 1 g salvia and 8 g mint
Summary of ScienceDome operations: Laia Ribas, Neus Sabaté and Cesca Cufí have been working on their experiments.
Dual split: Off
Summary of RAM operations: none
Summary of any observatory issues:
Robotic observatory: Not working. Replacement parts are scheduled to be installed during crew 261 rotation.
Summary of health and safety issues: none
Summary of other system’s issues: Some bridles have been added in the tunnels.
Questions, concerns and requests to Mission Support:
Spirit rover: replacement of the corroded wire – expected next week
Kitchen: knobs replacement – expected next week

Journalist Report – April 22nd

Crew 280 Journalist Report 22-04-2023

Author: Núria Jar, Crew Journalist

Saturday’s sun came without warning. Hypatia I crew members woke up lazily and they slowly sat around the table to have a coffee and weekend homemade pancakes. It was a peaceful and relaxing start for the Martian day SOL 6. Some were eager to continue with their experiments, while others were looking forward to exploring the outside. But all had one duty in mind: clean the station and its surroundings to maintain the habitability of Mars.

Three of the crew members conducted an extravehicular activity (EVA) around the MDRS to collect some bridles which were left on the ground. Hypatia I is very conscious about the importance of maintaining the integrity of the environment, especially to reduce the impact of human presence. The preservation of Mars and space is important for many reasons if we want to transform society into an interplanetary species.

Within the station, after a few hours of hard work, the Hab was looking pristine, ready for another week of experiments and exploration. The cleaning was reflected in the water consumption. From 11 liters per person on average, last day the consumption increased to 16 liters per person. Anyway, Hypatia I is proud of their water consumption during the first half of the simulation. They are learning to manage their daily routine with less water, which is crucial for long-duration space missions.

Another crucial routine for the crew is doing exercise in the morning. The Hab’s lower deck was transformed into a gym, where crew members can maintain their physical health. The back-up Crew Engineer, Helena Arias, developed personalized exercises tables for all crew members that can be done in a small space in a situation of confinement. Exercise is crucial for preventing muscle and bone loss, maintaining cardiovascular health, and improving overall well-being.

Finally, the crew flew a small drone around the base and took amazing pictures, such as one of the Crew Commander, Mariona Badenas-Agustí, from the window, where she said “Loud and clean.”

GreenHab Report – April 22nd

Crew 280 Green Hab Report 22-04-2023

GreenHab Officer: Laia Ribas

Environmental control: Open during the day and heater during the night

Temperatures: 80.5 F

Hours of supplemental light: N/A

Daily water usage for crops: 14 gallons

Daily water usage for research and/or other purposes: 0

Water in Blue Tank: 117 gallons

Time(s) of watering for crops: 08.00, 16:30 h

Changes to crops: N/A

The general watering of plants and crops was done. The door was opened in the morning due to high temperatures and closed during the afternoon due to the wind.

Harvest: 234 g cucumber, 112 g tomatoes, 4 g chives, 15 g jazz mix, 1 g salvia and 8 g mint

Support/supplies needed

EVA Report – April 22nd

Crew 280 EVA Report 22-04-2023

EVA #6

Author: Ariadna Farrés (HSO)

The EVA had two goals:
Drone flight around the base, with the purpose of continuing training drone maneuvers and to take some nice pictures/videos of the base for outreach purposes, by Carla Conejo (XO & Biologist) and Ariadna Farrés (Scientist & HSO).
Clean the plastic debris scattered around the base during the high-wind past days; Laia Ribas (GreenHab Officer).

Start time: 10:00 AM

End time: 11:45 AM

Narrative: After performing 5 minutes of depressurization in the front airlock, the EVA team started to walk towards Marble Ritual to find a good spot to start flying the drone to take some nice pictures/videos of the base. Three flights with the drone were performed: for the first flight the crew stayed in Marble Ritual and the drone flew towards the base and back where we were able to take some nice videos and pictures from the air. During the second flight the drone followed the crew back to the base, and the third flight was around the base.

Once the crew was back to base, the crew spent approximately 30 min walking around collecting plastic and metal debris that were scattered because of the winds of the last few days. The crew collected a total of 170g plastic debris, including bridles and other plastics, and 204g of metal debris, including wires and old nails.

Before entering the Hab, we followed the 5-minutes pressurization protocol.

Destination: Around the base and no further than Marble Ritual.

Coordinates (use UTM WSG 84): 518000 E, 4251000 N

EVA participants: 3: Ariadna Farrés (Scientist & HSO) (EVA Leader), Carla Conejo (XO & Biologist), Laia Ribas (GreenHab Officer).

Road(s) and routes per MDRS Map: Walk around the base and Cow Dung Road 0110

Mode of travel: Walk around the base and no further than 500m until Marble Ritual.

Copyright © The Mars Society. All rights reserved. | Main Site